OpenStudy (anonymous):

Jill starts to save at age 25 for a vacation home that she wants to buy for her 50th birthday. She will contribute $2500 each year to an account, which earns 1.3% annual interest, compounded quarterly. What is the future value of this investment, rounded to the nearest dollar, when Jill is ready to purchase the vacation home? $73,712 $73,951 $294,847 $295,806

OpenStudy (anonymous):

year one: 2500*1,013=x1 (x1+2500)*1,013=x2 (x2+2500)*1,013=x3 (....) (x25+2500)*1,013=answer AKA (2500*years)*(1.013^years)=which is none of the options. unless i misunderstood the question